OpenStudy (anonymous):

Give one (out of many) equations that is equivalent to 5x + 3y = 0

OpenStudy (anonymous):

i guss 5x-3y=0

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

10x+6y=0; 5x-3y would only be equvilant if y=0

OpenStudy (radar):

Sorry reply meant for another question

OpenStudy (anonymous):

any constant k multiplied by the equation if it is set equal to 0. Ex: 5/2x + 3/2y = 0